# Elucidating the representation of images within an unconditional diffusion model denoiser

### v1

@@ A.3 Training
Appendix B Stochastic reconstruction algorithm
We build our sampling algorithm based on the algorithm in [33] which does not require the noise level and follows an adaptive step size schedule. Hence, this network does not take the noise variance as an input and is a blind denoiser. This setup has the advantage that simplifies the network, which is important for analysis of the internal layers of the model.
Figure 10: More example samples. The samples are visually similar to the target image at the center of each panel. Interestingly, the location of the large, long-ranging image patterns are persevered in the samples, while the fine structures and details are diverse in their location with respect to the target image. See caption of Figure 3.
Figure 11: A block diagram illustrating the matching step implementation and its effect. Top: One forward pass through the UNet without matching guidance. Bottom: One forward pass after matching representations in the middle block. The noise initialization, $x_{t}$, are identical in both passes, but the output of the network is vastly different for them, due to the effect of matching. Matching helps the large structures in the conditioner image appear in the output image. See more examples of this effect in Figure 12.
Figure 12: Effect of matching on the output of the denoiser. Top row shows 4 noise samples (akin to initialization in the sampling trajectory). Leftmost coloum shows 4 conditioner target images. In the middle, each row shows the result of matching the $\phi$ of the noise initialization with the target’s and then one pass through the denoiser (score step). The rightmost column shows the output of the same network without matching (only the score step).
@@ Appendix B Stochastic reconstruction algorithm
Appendix C Properties of representation
C.1 Channel selectivity
Threshold.
To set the threshold, when computing the selectivity of channels, Figure 4, we compute the $\phi$’s at zero noise level for all dataset. In the ideal case, $\phi$ should be zero, but in practice some almost 1/3 of the channels are near-zero. This small amplitude can serve as a channel-wise threshold: For each channel, we set the threshold to be equal to maximum average activation for the channel at $\sigma=0$.
C.2 Clustering
We used python implementation of K-Means clustering algorithm, from sklearn package.
Clustering at different noise levels leads to similar results, consistent with stability of $\phi$ over noise levels shown in Figure 2. This is not true when noise is very small, since $\phi$ collapses to zero at small noise levels.
Figure 14: Continused from Figure 8. Random images from different clusters are shown in each row. Different cars with the same orientation are clustered together, but similar car brands with different orientations are assigned to different clusters. (see rows 2 and 3).

### v2

@@ A.3 Training
Appendix B Properties of representation
B.1 Representation sparsity
Figure 10: Change in sparsity for representations in models trained on Texture, LSUN-Bedroom, and CelebA datasets, from top to bottom row. Channel sparsity increases in the Middle block across dataset. See caption Figure˜1 for detailed description.
Figure 11: Stability of $\bar{a}$ across noise levels, for different network blocks trained on Texture, LSUN-Bedroom, and CelebA dataset. See caption of Figure˜2 for detailed description.
B.2 Channel selectivity
Figure 12: Channel selectivity for three models trained on Texture, LSUN-Bedroom, and CelebA datasets. Texture dataset: Distribution of channel selectivity shows that the vast majority of channels have low PR, corresponding to channels that are highly specialized (and infrequently active), concentrated on the left. The panels show the set of images that maximally activate each of four specialized channels, revealing selectivity for different texture patterns. PR of channels from left to right for texture model: $0.013,0.024,0.027,0.038$. LSUN-Bedroom: Channel selectivity is distributed more evenly with more common channels compared to the models trained on ImageNet and Texture datasets. This can be attributed to the presence of more common patterns and structures within the images of the dataset, as they all depict bedrooms. The PR of the four channels shown are $0.015,0.053,0.018,0.36,$, from left to right. CelebA: Most channels in this model respond to the majority of images. This can be attributed to the fact that all images in the dataset share the global layout and coarser level structures since the faces are aligned and centered. The four channels shown here have PR of $0.29,0.309,0.36,0.50$.
Figure 13: Channel selectivity predict signature statistical properties of the channels. From top to bottom row, models trained on Texture dataset, LSUN-bedroom, ImageNet and CelebA. From left to right: 1) marginal distribution of $\phi[i]$ in 4 specialized channels. 2) marginal distribution of $\phi[i]$ in 4 common channels. 3) channel selectivity is correlated with spatial sparsity within the channel. 4) The spatial variance of specialized channels is explained with fewer dimensions.
B.3 Semantic similarities in representation space
clustering.
We used the python implementation of K-Means clustering algorithm, from the sklearn package.
Clustering at different noise levels leads to similar results, consistent with stability of $\phi$ over noise levels shown in Figure˜2. This is not true when noise is very small, since $\phi$ collapses to zero at small noise levels.
Figure 15: Continued from Figure˜6. Random images from different clusters are shown in each row. Different cars with the same orientation are clustered together, but similar car brands with different orientations are assigned to different clusters. (see rows 2 and 3).
Figure 16: Random images from different clusters are shown in each row. Obtained from the model trained on Texture dataset.
Appendix C Stochastic reconstruction algorithm
We build our sampling algorithm based on the algorithm in [Kadkhodaie and Simoncelli, 2021] which does not require the noise level and follows an adaptive step size schedule. Hence, this network does not take the noise variance as an input and is a blind denoiser. This setup has the advantage that simplifies the network, which is important for analysis of the internal layers of the model.
Figure 17: More samples from model trained on Texture dataset. See caption of Figure˜7.
Figure 18: More samples from model trained on ImageNet64 dataset. The samples are visually similar to the target image at the center of each panel. Interestingly, the location of the large, long-ranging image patterns are persevered in the samples, while the fine structures and details are diverse in their location with respect to the target image. See caption of Figure˜7.
Figure 19: Samples generated unconditionally from models trained on Texture, CelebA, LSUN-Bedrooms, and ImageNet64 datasets respectively.
@@ Appendix C Stochastic reconstruction algorithm
Figure 21: A pair of clean images, $x$ and $x^{\prime}$, from ImageNet dataset is shown on the first column. The distance between the the two densities induced by the $\phi$’s obtained from these two images can be defined and computed using Equation˜4. The first term in the integrand is the difference between two conditional scores: the score of $x_{\sigma}$ given $\phi$ and the score of $x_{\sigma}$ given $\phi^{\prime}$. This difference is computed and integrated at all $\sigma$ levels (and symmetrized). For these two images, this difference is visualized in the green versus red boxes.